A gas cylinder of half argon and half helium has a total pressure of 1.11 tour, what is the partial pressure of each gas?
1
Expert's answer
2019-04-26T00:58:53-0400

Solution.

"p = p1 + p2 = R*T\/V*(n1+n2)"


1.11 tour = 147.99 Pa

p = RT/V*(1/2+1/2) = RT/V


"p1 = R*T*n1\/V"

p1 = 147.99*0.5 = 73.995 Pa


"p2 = p-p1"

p2 = 147.99-73.995 = 73.995 Pa

Answer:

p1 = 73.995 Pa

p2 = 73.995 Pa